OpenStudy (anonymous):

simplify and state excluded values. 5n-15/4n-12

OpenStudy (callisto):

\[\frac{5n-15}{4n-12} = \frac{5(n-3)}{4(n-3)}\] How would you simplify it?

OpenStudy (anonymous):

do u cros out n-3 and n-3

OpenStudy (callisto):

Yes. And what do you get for your answer?

OpenStudy (anonymous):

5/4 and what are the ecluded values

OpenStudy (callisto):

To find excluded value, put denominator =0, that is 4n-12 =0 Solve n.

OpenStudy (anonymous):